From: harvanek@apo.nmsu.edu Subject: 2.5-m Obslog 20010819 (MJD 52141) Date: Mon, 20 Aug 2001 06:33:49 -0600 (MDT) APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday August 19, 2001 (MJD 52141) ---=== OBSERVING TEAM ===--- Swing: Atsuko Nitta Kleinman (APO) Night: Mike Harvanek (APO) Pete Newman (APO) ---=== OBSERVING PLAN ===--- Science. ---=== OBSERVING SUMMARY ===--- 19:15 Doors open, fans on 01:15 Enclosure off 05:40 Enclosure on, doors closed, fans off Plate 651 completed. Good data for plate 663 (cartridge 3). ---=== OBSERVING LOG ===--- The dessicant on the east wall of the lower enclosure is mostly pink. The spectrograph chiller pressure remains at 36-38. The grade 5 N2 at the east wall of the lower enclosure dropped from 1800 to 1750 tonight so that leak may be back. Regarding the flaky humidity sensor last night, it was initially thought that this was just the sensor doing its calibration. However, Craig L. checked the data and said that it did not appear to be flagged as a calibration. Nor did the problem reoccur at 03:00 tonight as it should have if it was the calibration. The sensor appeared to behave normally tonight although the humidity was much lower than last night and so was not monitored as closely. Perhaps the problem was due to an insect in the sensor. We noticed that the blue dewar (#3) on sp2 is CAKED in ice whereas the other 3 dewars have little or no ice on them. Perhaps this is related to the "dewar 3 ln2 empty" watcher error (mentioned below) and the related critical error false alarms that occurred around 02:00. There was a high pitched whistling coming from the vent line in the plastic cap of this same dewar when we closed. Neither of us could recall hearing that sound before. Has anyone checked to see if the telescope can collide with the new weather station located on the SW corner of the platform? We looked at this tonight by bringing the telescope down to an altitude of 20 near the station. It appears that it's possible for the telescope to hit the station and the flat field petal will certainly hit it. Are there going to be new altitude limits for azimuths outside the 120.8 - 121.4 range? Has this been taken into consideration for the DIMM mounting? The humidity stayed low tonight but clouds kept us from removing the enclosure until 01:15. After we got on the sky we experienced a few problems. StartGuider failed after one iteration with this in murmur Aug 20 01:45:47 sdsshost IOP 143335 TEXTONLY expected floating-point number but got "NaN" Restarting the guider solved the problem. Also, it seems that we are running with the 20-star pointing model obtained on (2001/8/12, MJD52134, tdat:telmod.dat;21) no doubt reloaded a few days ago when telstop/telrun commands were issued. The calibration offsets reflect this: CalAzOff = 0.003644 = 13 arcsec CalAltOff= -0.000029 = -0.1 arcsec CalAngOff= -0.025329 = 9 arcsec However, we are finding guide stars on the first attempt (3 plates in a row now, dating back to the previous night) so we are leaving well alone for now. Given the known problems with this telescope model (12arcsec systematic offset to east), it is suspected that celestial-northern fields will NOT find stars, and we will revert to the preceeding model then (telmod.dat;20), this time copying it to a new file to make it the latest version (hence telmod.dat;22) so it gets loaded by default next time. The calibration offsets should then be cleared. Plate 651, cartridge 5: preCalibs aborted with an error window saying: "Error: shSubRegNew: subregion row/col out of bounds" on the gflat analysis, switching startMon from plate-specific to simple mode. The traceback was lost due to a typo. A preCalibs command from the command-line interface worked. At the same time, watchers reported a ln2 fill fault on sp2. It was noted that during a preCalib sequence, the flat-field lamps remain on well until after the exposure reads out. This could be changed to increase lamp lifetime (although no observing efficiency is lost). It was also noted that the arcs seemed to be on during the spectro flats. Is that correct?! Also got this, once: Mon Aug 20 02:16:58 MDT 2001: Spectro sp1 mechanicals parse error: n empty or garbled reply from sp1 mechanicals We also continued to see the following error (many times): Modu="mir o MoveOneGalil" (Aug 20 02:20:52 MDT) ---=== SPECTROSCOPY DATA SUMMARY ===--- UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 10584, plate -9999 ------- 23:22 10584 0.0 bias {{}} 23:26 10585 0.0 bias {{}} 00:42 10586 0.0 bias {{}} ----- sequence 10587, plate 651 ------- 16.7 15.2 13.9 12.9 08:09 10587 10.0 flat pre-flat (declared done) 08:11 10588 2.0 arc pre-arc 08:34 10589 901.0 target science 08:52 10590 901.0 target science 09:14 10591 902.0 target science 09:35 10592 1052.0 target science 09:42 10593 240.0 smear {{Smear Pattern}} 09:47 10594 10.0 flat post-flat 09:49 10595 2.0 arc post-arc ----- sequence 10596, plate 663 ------- 12.8 11.4 10.1 8.3 10:09 10596 10.0 flat pre-flat 10:11 10597 2.0 arc pre-arc 10:30 10598 900.0 target science 10:48 10599 901.0 target science 11:15 10600 1502.0 target science 11:18 10601 10.0 flat post-flat 11:20 10602 2.0 arc post-arc Plate 651 was observed in sequence(s) 10587 for a total of 3756 seconds Plate 663 was observed in sequence(s) 10596 for a total of 3303 seconds ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Cart Plate getclock Az Alt Rot Scale ----------------------------------------------------------------- 02:18 5 651 998295495 0.003644 -0.000029 -0.025329 0.999950 04:21 5 651 998302865 0.003087 -0.001450 0.027489 0.999900 ---=== DATA TAPE SUMMARY ===--- Ship to FNAL: JL4293 Stays at APO: JL4294 ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm " -------------------------------------------------------------------------------- 19:15 Doors open, fans on 01:15 Enclosure off 02:18 5 651 0.99995 493 1397 -210 21 45.0 14.2 10 133 - 1.7 02:45 5 651 0.99995 493 1326 -275 11 46.6 14.2 14 142 - 2.0 02:47 5 651 0.99995 493 1376 -225 11 46.7 14.2 17 138 - 1.7 02:59 5 651 0.99995 493 1449 -150 6 47.0 14.0 15 144 - 1.8 04:21 5 651 0.99990 986 1877 -130 10 47.3 13.2 16 164 - 1.8 05:40 Enclosure on, doors closed, fans off ---=== WEATHER LOG ===--- Variable clouds throughout the night. Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky Precip. ------------------------------------------------------------------- 16:01 64 48 5 249 (WSW) 958 - pcl none 17:01 65 46 0 245 (WSW) 908 - pcl none 18:01 65 45 0 232 (SW) 939 - pcl none 19:01 64 45 1 358 (N) 929 - pcl none 20:01 63 45 1 136 (SE) 942 - mcl none 21:01 63 30 10 123 (ESE) 356 - pcl none 22:01 62 32 9 159 (SSE) 571 - pcl none 23:01 60 34 16 144 (SE) 842 - pcl none 00:01 59 35 14 138 (SE) 978 - pcl none 01:01 59 35 11 148 (SSE) 995 - pcl none 02:01 58 34 13 137 (SE) 1029 - pcl none 03:01 57 31 15 143 (SE) 933 - pcl none 04:01 56 36 17 160 (SSE) 1477 - pcl none 05:01 55 41 12 181 (S) 1261 - pcl none [pmt]cl = partly, mostly, totally cloudy ---=== LN2 SCHEDULE ===--- Spectrographs filled at 01:20. Spectro dewar weight at 06:00: 202 lb. Imager dewar weight at 06:00: 200 lb. ---=== TELESCOPE STATUS ===--- Stowed at 30 deg alt with cartridge 3 (plate 663) mounted. Counterweights at 230. Spectro auto-fill connected. ---=== SOFTWARE USED ===--- IOP/SOP: v3_53_6 Watcher: v2_9_0 MCP: v5_3_1 TPM: v2_4_1 AstroDa: v14_36 TCC: 2.5.5 sdssProcedures: v1_22 tccMon : v1_6 ---=== MIRROR NUMBERS ===--- Time dpTemp dewPt Press TmpOut HumOut Windd Winds 21:56 61.8 32.0 21.68 32 141.6 12.0 16 PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A: 0.0720 0.0710 Axial B: -0.0880 -0.0980 Axial C: 0.0650 0.0610 Trans D: 4.1850 4.1800 Lateral E: 1.9335 1.9365 Lateral F: 0.4340 0.4370 GALILS Commanded: 5400. -3700. 900. -200. 31550. 30650. Actual: 5400. -3700. 900. -200. 31550. 30650. SETMIR VALUES primDesOrient: 0.00 -11.80 23.00 1256.90 642.10 primOrient: 0.00 -12.16 22.81 1257.41 642.19 SECONDARY: ---------- Focus: 0.0 Air temp: 16.6 C Altitude: 30.41 MIGS TONIGHT NOMINAL Axial A: -1.2940 -1.2900 Axial B: -1.3030 -1.3150 Axial C: -1.3130 -1.3060 Trans D: 0.1350 0.1890 GALILS Commanded: 1583850. 1583850. 1583850. 2600. 3800. Actual: 1583832. 1583838. 1583882. 2600. 3800. SETMIR VALUES secDesOrient: 1257.00 0.00 0.00 0.00 -175.00 secOrient: 1257.01 -0.02 0.01 0.00 -175.11 ---=== PROBLEMS IN DETAIL ===--- No details.